In planning for a new item, a manufacturer assumes that the number of items produced x and the cost in dollars C of producing these items are related by a linear equation.Projections are that 100 items will cost 10,000 to produce and that 300 items will cost 22,000 to produce. find the equation that relates C and x
Answer by stanbon(48535)

Projections are that 100 items will cost 10,000 to produce and that 300 items will cost 22,000 to produce. find the equation that relates C and x
You have two points: (100,10000) and (300,22000)
The equation you want has the form C=mx+b
m=[22000-10000]/[300-100]=12000/200=60
With Cost=10000, x=100 and m=60 you can find "b", as follows:
10000=60(100)+b
10000=6000+b
b=4000
EQUATION:
C=60x+4000
Cheers,
stan H.